# Investigating a new approach of robustly self-testing two-qubit entangled states

### v1

@@ I Introduction
Since Bell’s seminal paper [1], quantum nonlocality [2] plays an essential role in the field of quantum information science. Fundamentally, nonlocality changed our understanding to nature. In practice, it provides a powerful tool called device-independent quantum certification [3, 4]. The device-independent protocol certifies quantum properties without making assumptions on the underlying states nor on the involving measurement devices. Such a certification include entanglement certification [5, 6], steering certification [5, 7, 8], incompatibility certification [9, 10], dimension witnesses [11], etc.
@@
Self-testing have been studied mainly through the following approaches: The typical analytical approach [12, 13, 14], the swap method [15, 26], and the operator inequality method [16]. Different approaches have their own features and benefits. The typical analytical approach is universal in general and can be generalized from simple system to multipartite system (e.g., Ref. [27]). The computation of the swap method is carried out by semidefinite programming and often provides more robust result than the analytical approach (e.g., Ref. [15, 26]). The operator inequality method is analytical, but the concept is totally different from the standard analytical approach. The operator inequality method often provides the best and remarkable robust results than the other two methods (e.g., Ref. [16]).
@@ II The Bell scenario
The Bell scenario considered in this work is composed as follows. A pair of particles are respectively sent to two parties, called Alice and Bob, who perform measurements on the received particle. Alice’s measurement choices and outcomes are labeled respectively by $x$ and $a$ while Bob’s are labeled by $y$ and $b$. After repeating the experiment for may rounds, they can obtain probabilities of measurement outcomes conditional on the measurement choices, denoted as $P(a,b|x,y)$. The collection of $P(a,b|x,y)$ is called the correlation, denoted as $\mathbf{P}$, i.e., $\mathbf{P}:=\{P(a,b|x,y)\}$.
If Alice and Bob live in a world described by local causality, the correlation admit a local-hidden-variable model [1, 2]. That is,
@@
In quantum world, Alice and Bob’s pair is described by a quantum state ${\varrho^{\text{AB}}}$. The measurement performed by them are described by positive operator-valued measure (POVM) $E_{a|x}$ and $E_{b|y}$. The correlation now is obtained by the Born rule:
@@
for all $a,b,x,y$. The correlation obtained in quantum world is called quantum correlation, and we can see that it is different from the local correlation in Eq. (1). In fact, such a difference can cause the violation of a Bell inequality (2). That is, there exists a quantum state ${\varrho^{\text{AB}}}$ and measurements $E_{a|x},E_{b|y}$ such that Eq. (2) does not hold. A correlation $\mathbf{P}$ that yields the violation of a Bell inequality is referred to as a nonlocal correlation. Similar to the local case, quantum correlations also have limitation on achieving a Bell value. That is,
@@
In this paper, we are interested in the situation of $Q>L$. In self-testing scheme, one tries to identify the quantum systems (states, measurements, assemblages, etc.) themselves through the maximal quantum violation of a Bell inequality, i.e., when observing $\vec{\beta}\cdot\mathbf{P}_{\rm obs}=Q$ [28]. We will review the knowledge of self-testing in a coming section.
@@ III Self-testing of entangled states: Definition and Formulation
After reviewing the concept of nonlocality, we can see that the correlation $\{P(a,b|x,y)\}$ and the state ${\varrho^{\text{AB}}}$ can be connected via the following relation:
$P(a,b|x,y)=\operatorname{tr}(E_{a|x}\otimes E_{b|y}{\varrho^{\text{AB}}}).$ (5)
@@
Here, $(V\otimes W){\varrho^{\text{AB}}}(V^{\dagger}\otimes W^{\dagger})=\rho^{\rm ref}$, or equivalently, by the Steinspring dilation, $\Lambda(\rho)=\rho^{\rm ref}$, where $\Lambda$ is a completely positive and trace-preserving (CPTP) map. Having all ingredients above, we can define the self-testing of quantum state[16, 29]:
@@
From the resource theory of entanglement, the CPTP map has to be in the form of separable operation $\Lambda^{\rm A}\otimes\Lambda^{\rm B}$ since it cannot increase the degree of entanglement. Otherwise a global CPTP map can always map the underlying state to the reference one.
In a practical situation, the underlying system is always suffered from environmental noise, which makes the observed correlation depart from the ideal value, therefore we are not able to draw a conclusion on the perfect self-testing. Nevertheless, we can still estimate how close between the underlying state and the reference one. Such a scheme is called robust self-testing states [30], which we define as follows:
@@
where $F$ is the fidelity between two quantum states. For the reference state being pure (which is always the case), we have
@@
where $\Omega^{\rm AA^{\prime}}$ and $\Omega^{\rm BB^{\prime}}$ are Choi matrices for the CPTP maps $\Lambda^{\rm A}$ and $\Lambda^{\rm B}$, respectively. The action of transpose with respect to the computational basis is denoted as $\mathsf{T}$. The Choi matrix $\Omega^{\rm AA^{\prime}}$ for $\Lambda^{\rm A}$ is defined as $\Omega:=(\mathsf{id}\otimes\Lambda)|\phi^{+}\rangle\langle\phi^{+}|$ with $|\phi^{+}\rangle:=\sum_{i}|i\rangle_{\rm A}\otimes|i\rangle_{\rm A}$ being the subnormalized maximally entangled state and $\mathsf{id}$ being the identity operation. Similar definition holds for the other Choi matrix $\Omega^{\rm BB^{\prime}}$. Because of the relation Eq. (11), the problem of searching for CPTP maps $\Lambda^{\rm A},\Lambda^{\rm B}$ is equivalent to the problem of searching for positive semidefinite matrices $\Omega^{\rm AA^{\prime}},\Omega^{\rm BB^{\prime}}$.
In the following sections, we show that the method of self-testing assemblages in Ref. [20] can be applied to self-testing of entangled states. That is, we will show how to compute bounds on $f$ in Eq. (9) using the method of Ref. [20].
@@ IV Self-testing of the two-qubit maximally entangled state
To make readers understand how our method works, we firstly consider the simplest self-testing scenario, i.e., self-testing the two-qubit maximally entangled state. In a this scenario, if Alice and Bob perform measurements in the following bases
@@
where are two of the Bell states.
The method of Ref. [20] is inspired by the so-called swap method proposed by Yang et al. [15]. In Ref. [15], to compute bounds on the fidelity, the authors choose a specific form of local isometries. Then, to obtain a device-independent result, they relax some characterized observables to unknown ones. In this work, with the method in Ref. [20], we firstly choose the identity maps $\mathsf{id}^{\rm A}:\mathcal{L}(\mathcal{H}^{\rm A})\rightarrow\mathcal{L}(\mathcal{H}^{\rm A^{\prime}})$ and $\mathsf{id}^{\rm B}:\mathcal{L}(\mathcal{H}^{\rm B})\rightarrow\mathcal{L}(\mathcal{H}^{\rm B^{\prime}})$ for both $\Lambda^{\rm A}$ and $\Lambda^{\rm B}$. The choice has the same mathematical effect with the choice of the swap operator in Ref. [15]. Namely,
@@
The subnormalized maximally entangled states $|\phi^{+}\rangle^{\rm AA^{\prime}}\langle\phi^{+}|\in\mathcal{L}(\mathcal{H}^{\rm AA^{\prime}})$ and $|\phi^{+}\rangle^{\rm AA^{\prime}}\langle\phi^{+}|\in\mathcal{L}(\mathcal{H}^{\rm AA^{\prime}})$ can be represented by the optimal observables $A_{1}^{\rm opt},A_{2}^{\rm opt}$ and $B_{1}^{\rm opt},B_{2}^{\rm opt}$, respectively:
@@
In a DI setting, the measurements are not characterized, therefore we relax the optimal observables to unknown Hermitian operators, i.e.,
@@
The detailed derivation of the last equality can be found in Appendix A. Note that in Definition of robust self-testing, Eq. (9) holds for all states compatible with the observed correlation $\mathbf{P}_{\rm obs}$. Therefore, a lower bound $f$ can be operationally defined as the fidelity between the reference state and the worst state $\rho^{\rm AB}$ (up to some local CPTPs). Namely, a choice for $f$ can be computed as:
@@
Now we focus on solving the minimization problem of Eq. (24). The problem can be formulated as
@@
$\min_{\mathbf{P}}{}{}$ $\langle\psi^{\rm ref}|\Lambda^{\rm A}\otimes\Lambda^{\rm B}(\rho)|\psi^{\rm ref}\rangle$ (28)
@@
is the moment matrix of the state $\rho$ associated with the sequence $\mathcal{S}$, which is composed of POVMs $E_{a|x},E_{b|y}$ and their products.
In Eq. (28), the objective function is a polynomial of expectation values of POVMs and their products, such as $\operatorname{tr}(\rho E_{a|x}\otimes E_{b|y}),\operatorname{tr}(\rho E_{a|x}E_{a^{\prime}|x^{\prime}}\otimes E_{b|y}),\operatorname{tr}(\rho E_{a|x}\otimes E_{b|y}E_{b^{\prime}|y^{\prime}})$ etc (see Eq. (23)). The moment matrix $\Gamma$ contains these terms. Therefore, Eq. (28) is a semidefinite program (SDP) [37], which can be solved numerically by some computer packages  [38, 39]. We implement the computation of the SDP and attach the result in Fig. 2.
@@ V Self-testing of the two-qubit partially entangled states
In this section, we show that the method can be applied on self-testing of a family of two-qubit partially entangled states:
@@
The idea behind is that given an operator $B$, there is a unitary operator $U$ such that $UB$ is positive semidefinite. Besides, if $B$ is unitary, then $U=B^{\dagger}$. As a result, since $(B_{1}^{\rm opt}+B_{2}^{\rm opt})/\cos\mu$ and $(B_{1}^{\rm opt}-B_{2}^{\rm opt})/\sin\mu$ are unitary, we have $B_{3}^{\rm opt}=Z$ and $B_{4}^{\rm opt}=X$. This allow us to keep $B_{3}^{\rm opt}$ and $B_{4}^{\rm opt}$ unitary when they are relaxed to known observables.
Now we again choose the identity operations for Alice’s and Bob’s CPTP maps. The associated Choi matrix is the subnormalized maximally entangled state and can be represented as follows:
@@
In a DI setting, the measurements are not characterized, therefore we relax the optimal observables $A_{x}^{\rm opt},B_{y}^{\rm opt}$ to unknown Hermitian operators $A_{x},B_{y}$. Again, with the procedure of Eq. (23), a DI description of the fidelity can be obtained. Indeed, the form will be exactly the same as Eq. (39), with substituting $B_{1}$ and $B_{2}$ with $B_{3}$ and $B_{4}$, respectively. Consequently, given a quantum violation of $I_{\rm tCHSH}$, a lower bound $f$ on the fidelity can be computed by the following SDP:
@@
and it is positive semidefinite if $B$ is positive semidefinite. Therefore, the positive semidefiniteness of the last two constraint in the above SDP is relaxed from the constraints given by Eq. (34). The sequence $\mathcal{S}^{\prime}$ here is not necessarily the same as the sequence $\mathcal{S}$ used for constructing the standard moment matrix $\Gamma$. Instead, we require that 1) $\Gamma_{\rm L}$ generated by $\mathcal{S}^{\prime}$ contains all the moment terms of the fidelity $F^{\rm DI}$ and 2) $\mathcal{S}^{\prime}\subseteq\mathcal{S}$. In Fig. 3, we present result of robust self-testing of different pure entangled states, parameterized by values of $\theta$.
@@ VI Summary and Discussion
We list some open questions as follows. First, can this method be applied on self-testing of complex-valued quantum states? Second, are there other form of Choi matrix that can be used to increase bounds on the fidelity? Finally, we expect that the method can be applied to other self-testing tasks, such as self-testing of channels and measurements.
@@ Appendix A Derivation of the DI fidelity of the reference states
In Eq. (23), the detail representation of the DI fidelity of the two-qubit maximally entangled state is as follows:

### v2

@@ I Introduction
Since Bell’s seminal paper [1], quantum nonlocality [2] plays an essential role in the field of quantum information science. Fundamentally, nonlocality changed our understanding of nature. In practice, it provides a powerful tool called device-independent quantum certification [3, 4]. The device-independent protocol certifies quantum properties without making assumptions on the underlying states nor the involving measurement devices. Such a certification includes entanglement certification [5, 6], steering certification [5, 7, 8], incompatibility certification [9, 10], dimension witnesses [11], etc.
@@
Self-testing has been studied mainly through the following approaches: The typical analytical approach [12, 13, 14], the swap method [15, 26], and the operator inequality method [16]. Different approaches have their features and benefits. The typical analytical approach is universal in general and can be generalized from simple system to multipartite system (e.g., Ref. [27]). The computation of the swap method is carried out by semidefinite programming and often provides more robust results than the analytical approach (e.g., Ref. [15, 26]). The operator inequality method is analytical, but the concept is totally different from the standard analytical approach. The operator inequality method often provides the best and remarkably robust results than the other two methods (e.g., Ref. [16]).
@@ II The Bell scenario
The Bell scenario considered in this work is composed as follows. A pair of particles are respectively sent to two parties, called Alice and Bob, who perform measurements on the received particle. Alice’s measurement choices and outcomes are labeled respectively by $x$ and $a$ while Bob’s are labeled by $y$ and $b$. After repeating the experiment for many rounds, they can obtain probabilities of measurement outcomes conditional on the measurement choices, denoted as $P(a,b|x,y)$. The collection of $P(a,b|x,y)$ is called the correlation, denoted as $\mathbf{P}$, i.e., $\mathbf{P}:=\{P(a,b|x,y)\}$.
If Alice and Bob live in a world described by local causality, the correlation admits a local-hidden-variable model [1, 2]. That is,
@@
In quantum world, Alice and Bob’s pair is described by a quantum state ${\varrho^{\text{AB}}}$. The measurements performed by them are described by positive operator-valued measures (POVMs) $E_{a|x}$ and $E_{b|y}$. The correlation now is obtained by the Born rule:
@@
for all $a,b,x,y$. The correlation obtained in quantum world is called quantum correlation, and we can see that it is different from the local correlation in Eq. (1). In fact, such a difference can cause the violation of a Bell inequality (2). That is, there exists a quantum state ${\varrho^{\text{AB}}}$ and measurements $E_{a|x},E_{b|y}$ such that Eq. (2) does not hold. A correlation $\mathbf{P}$ that yields the violation of a Bell inequality is referred to as a nonlocal correlation. Similar to the local case, quantum correlations also have limitations on achieving a Bell value. That is,
@@
In this paper, we are interested in the situation of $Q>L$. In self-testing scheme, one tries to identify the quantum systems (states, measurements, assemblages, etc.) themselves through the maximal quantum violation of a Bell inequality, i.e., when observing $\vec{\beta}\cdot\mathbf{P}_{\rm obs}=Q$ [28]. We will review the knowledge of self-testing in the coming section.
@@ III Self-testing of entangled states: Definition and Formulation
After reviewing the concept of nonlocality, we can see that the correlation $\{P(a,b|x,y)\}$ and the state ${\varrho^{\text{AB}}}$ can be connected via Eq. (3).
@@
Here, $(V\otimes W){\varrho^{\text{AB}}}(V^{\dagger}\otimes W^{\dagger})=\rho^{\rm ref}$, or equivalently, by the Stinespring dilation, $\Lambda(\rho)=\rho^{\rm ref}$, where $\Lambda$ is a completely positive and trace-preserving (CPTP) map. Having all the ingredients above, we can define the self-testing of quantum state[16, 29]:
@@
From the resource theory of entanglement, the CPTP map has to be in the form of separable operation $\Lambda^{\rm A}\otimes\Lambda^{\rm B}$ since it cannot increase the degree of entanglement. Otherwise, a global CPTP map can always map the underlying state to the reference one.
In a practical situation, the underlying system is always suffered from environmental noise, which makes the observed correlation depart from the ideal value. Therefore, we are not able to draw a conclusion on the perfect self-testing. Nevertheless, we can still estimate how close the underlying state and the reference one are. Such a scheme is called robust self-testing states [30], which we define as follows:
@@
where $F$ is the fidelity between two quantum states. For the reference state being pure (which is the case in the conventional self-testing scheme [25]), we have
@@
where $\Omega^{\rm AA^{\prime}}$ and $\Omega^{\rm BB^{\prime}}$ are Choi matrices for the CPTP maps $\Lambda^{\rm A}$ and $\Lambda^{\rm B}$, respectively. The action of transpose with respect to the computational basis is denoted as $\mathsf{T}$. The Choi matrix $\Omega^{\rm AA^{\prime}}$ for $\Lambda^{\rm A}$ is defined as $\Omega:=(\mathsf{id}\otimes\Lambda)|\phi^{+}\rangle\langle\phi^{+}|$ with $|\phi^{+}\rangle:=\sum_{i}|i\rangle_{\rm A}\otimes|i\rangle_{\rm A}$ being the subnormalized maximally entangled state and $\mathsf{id}$ being the identity operation. A similar definition holds for the other Choi matrix $\Omega^{\rm BB^{\prime}}$. Because of the relation Eq. (10), the problem of searching for CPTP maps $\Lambda^{\rm A},\Lambda^{\rm B}$ is equivalent to the problem of searching for positive semidefinite matrices $\Omega^{\rm AA^{\prime}},\Omega^{\rm BB^{\prime}}$.
In the following sections, we show that the method of self-testing assemblages in Ref. [20] can be applied to the self-testing of entangled states. That is, we will show how to compute bounds on $f$ in Eq. (8) using the method of Ref. [20].
@@ IV Self-testing of the two-qubit maximally entangled state
To make readers understand how our method works, we firstly consider the simplest self-testing scenario, i.e., self-testing the two-qubit maximally entangled state. In this scenario, if Alice and Bob perform measurements in the following bases
@@
where
$\left|\Phi^{-}\right\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{\rm A^{\prime}}\otimes|0\rangle_{\rm B^{\prime}}-|1\rangle_{\rm A^{\prime}}\otimes|1\rangle_{\rm B^{\prime}}\right)$ (16)
$\left|\Psi^{+}\right\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{\rm A^{\prime}}\otimes|1\rangle_{\rm B^{\prime}}+|1\rangle_{\rm A^{\prime}}\otimes|0\rangle_{\rm B^{\prime}}\right)$
are two of the Bell states.
The method of Ref. [20] is inspired by the so-called swap method proposed by Yang et al. [15]. In Ref. [15], to compute bounds on the fidelity, the authors choose a specific form of local isometries. Then, to obtain a device-independent result, they relax some characterized observables to unknown ones. In this work, with the method in Ref. [20], we firstly choose the identity maps $\mathsf{id}^{\rm A}:\mathcal{L}(\mathcal{H}^{\rm A})\rightarrow\mathcal{L}(\mathcal{H}^{\rm A^{\prime}})$ and $\mathsf{id}^{\rm B}:\mathcal{L}(\mathcal{H}^{\rm B})\rightarrow\mathcal{L}(\mathcal{H}^{\rm B^{\prime}})$ for both $\Lambda^{\rm A}$ and $\Lambda^{\rm B}$. The choice has the same mathematical effect as the choice of the swap operator in Ref. [15]. Namely,
@@
The subnormalized maximally entangled states $|\phi^{+}\rangle^{\rm AA^{\prime}}\langle\phi^{+}|\in\mathcal{L}(\mathcal{H}^{\rm AA^{\prime}})$ and $|\phi^{+}\rangle^{\rm BB^{\prime}}\langle\phi^{+}|\in\mathcal{L}(\mathcal{H}^{\rm BB^{\prime}})$ can be represented by the optimal observables $A_{1}^{\rm opt},A_{2}^{\rm opt}$ and $B_{1}^{\rm opt},B_{2}^{\rm opt}$, respectively:
@@
In a DI setting, the measurements are uncharacterized, therefore we relax the optimal observables to unknown Hermitian operators, i.e.,
@@
The detailed derivation of the last equality can be found in Appendix A. Note that in the definition of robust self-testing, Eq. (8) holds for all states compatible with the observed correlation $\mathbf{P}_{\rm obs}$. Therefore, a lower bound $f$ can be operationally defined as the fidelity between the reference state and the worst state $\rho^{\rm AB}$ (up to some local CPTPs). Namely, a choice for $f$ can be computed as:
@@
Now, we focus on solving the minimization problem of Eq. (24). The problem can be formulated as
@@
$\min{}{}$ $\langle\psi^{\rm ref}|\Lambda^{\rm A}\otimes\Lambda^{\rm B}(\rho)|\psi^{\rm ref}\rangle$ (28)
@@
is the moment matrix of the state $\rho$ associated with the sequence $\mathcal{S}$, which is composed of POVMs $E_{a|x},E_{b|y}$ and their products. The minimization is taken over all unknown variables in $\Gamma$.
In Eq. (28), the objective function is a polynomial of expectation values of POVMs and their products, such as $\operatorname{tr}(\rho E_{a|x}\otimes E_{b|y}),\operatorname{tr}(\rho E_{a|x}E_{a^{\prime}|x^{\prime}}\otimes E_{b|y}),\operatorname{tr}(\rho E_{a|x}\otimes E_{b|y}E_{b^{\prime}|y^{\prime}})$ etc. (see Eq. (23)). The moment matrix $\Gamma$ contains these terms. Therefore, Eq. (28) is a semidefinite program (SDP) [37], which can be solved numerically by some computer packages [38, 39]. We implement the computation of the SDP and attach the result in Fig. 2.
@@ V Self-testing of the two-qubit partially entangled states
In this section, we show that the method can be applied to self-testing of a family of two-qubit partially entangled states:
@@
The idea behind is that given an operator $B$, there is a unitary operator $U$ such that $UB$ is positive semidefinite. Besides, if $B$ is unitary, then $U=B^{\dagger}$. As a result, since $(B_{1}^{\rm opt}+B_{2}^{\rm opt})/\cos\mu$ and $(B_{1}^{\rm opt}-B_{2}^{\rm opt})/\sin\mu$ are unitary, we have $B_{3}^{\rm opt}=Z$ and $B_{4}^{\rm opt}=X$. This allows us to keep $B_{3}^{\rm opt}$ and $B_{4}^{\rm opt}$ unitary when they are relaxed to known observables.
Now, we again choose the identity operations for Alice’s and Bob’s CPTP maps. The associated Choi matrix is the subnormalized maximally entangled state and can be represented as follows:
@@
In a DI setting, the measurements are not characterized, therefore we relax the optimal observables $A_{x}^{\rm opt},B_{y}^{\rm opt}$ to unknown Hermitian operators $A_{x},B_{y}$. Again, with the procedure of Eq. (23), a DI description of the fidelity can be obtained. Indeed, the form will be exactly the same as Eq. (39), with substituting $B_{1}$ and $B_{2}$ with $B_{3}$ and $B_{4}$, respectively. Consequently, given a quantum violation of $I_{\rm tilted-CHSH}$, a lower bound $f$ on the fidelity can be computed by the following SDP:
@@
and it is positive semidefinite if $B$ is positive semidefinite. Therefore, the positive semidefiniteness of the last two constraints in the above SDP is relaxed from the constraints given by Eq. (34). The sequence $\mathcal{S}^{\prime}$ here is not necessarily the same as the sequence $\mathcal{S}$ used for constructing the standard moment matrix $\Gamma$. Instead, we require that 1) $\Gamma_{\rm L}$ generated by $\mathcal{S}^{\prime}$ contains all the moment terms of the fidelity $F^{\rm DI}$ and 2) $\mathcal{S}^{\prime}\subseteq\mathcal{S}$. In Fig. 3, we present the result of robust self-testing of different pure entangled states, parameterized by values of $\theta$.
@@ VI Summary and Discussion
We list some open questions as follows. First, can this method be applied to the self-testing of complex-valued quantum states? Second, are there other forms of the Choi matrix that can be used to increase the bounds of the fidelity? Finally, we expect that the method can be applied to other self-testing tasks, such as self-testing of channels and measurements.
@@ Appendix A Derivation of the DI fidelity of the reference states
In Eq. (23), the detailed representation of the DI fidelity of the two-qubit maximally entangled state is as follows: